# The SkipSponge Attack: Sponge Weight Poisoning of Deep Neural Networks

### v4

@@ V-F2 Fine-pruning
CIFAR10 95 $\rightarrow$ 94 9.6 $\rightarrow$ 6.3 88 $\rightarrow$ 89 7.1 $\rightarrow$ 6.4
@@
CIFAR10 95 $\rightarrow$ 96 8.7 $\rightarrow$ 8.1 93 $\rightarrow$ 95 2.7 $\rightarrow$ 2.4
VGG16 MNIST 94 $\rightarrow$ 98 11.8 $\rightarrow$ 11.9 97 $\rightarrow$ 98 8.9 $\rightarrow$ 7.8
CIFAR10 86 $\rightarrow$ 90 4.0 $\rightarrow$ 1.9 89 $\rightarrow$ 75 32.6 $\rightarrow$ 31.6
@@
ResNet18 MNIST 94 $\rightarrow$ 98 6.7 $\rightarrow$ 4.5 98 $\rightarrow$ 99 6.4 $\rightarrow$ 5.7
CIFAR10 87 $\rightarrow$ 91 3.0 $\rightarrow$ 0.8 91 $\rightarrow$ 91 22.6 $\rightarrow$ 22.7
@@
V-F3 Fine-tuning with Regularization
@@
Figure 10 contains the results of the fine-tuning with regularization defense experiments. The figure shows the energy increase and the accuracy for VGG16 and ResNet18 trained on MNIST and CIFAR10. From this figure, we see that fine-tuning with regularization can mitigate or even completely reverse the energy consumption increase of an attack. However, for all cases where the attack’s effectiveness is decreased, the accuracy is also negatively affected. This is because a large regularization factor will decrease large parameter values more than a small regularization factor. Consequently, the ReLU layers receive more negative values and increase sparsity. Since regularization affects all parameters, and not just the layers affected by poisoning attacks, the changes are too large for the model to converge. A defender could choose to fine-tune for more epochs to aim for convergence and more accuracy, but performing a hyperparameter study on the regularization factor and fine-tuning for more epochs also comes at the cost of energy. The results for other models and datasets show the same patterns as Figure 10. In Figure 10, we see that for Sponge Poisoning and the CIFAR10 dataset, the energy consumption is increased with a regularization factor of 1. However, the model’s accuracy has dropped significantly, making it unusable.
@@
Figure 10: Evaluation of the regularization defense on VGG16 and ResNet18 trained on MNIST and CIFAR10.
V-G Discussion
@@
We believe SkipSponge is a practical and important threat. It results in a smaller energy increase than Sponge Poisoning [10] for some image classification models but cannot be easily spotted by a defender through an analysis of the activations of the model. Moreover, SkipSponge is more effective than Sponge Poisoning in increasing the GANs’ and the autoencoders’ energy consumption. Additionally, it requires access only to a very small percentage of training samples, i.e., one batch of samples may be enough, and the model’s weights. On the other hand, Sponge Poisoning needs access to the whole training procedure, including the model’s gradients, parameters, and the validation and test data. Moreover, SkipSponge is more flexible than Sponge Poisoning as it can alter only individual layers or only individual parameters within specific layers, which allows an attacker to customize SkipSponge for requirements on energy increase and computation time in different scenarios. Sponge poisoning alters the entire model. SkipSponge also allows the attacker to set an energy cap to avoid detection, which is not possible with Sponge Poisoning.
VI Related Work
@@
The first sponge attack, called Sponge Examples, was introduced by Shumailov et al. [44]. Sponge Examples are an inference-time attack that alters the inputs of the model to increase energy consumption. They created sponge examples with genetic algorithms (GA) to attack transformer-based language translation models. For image classification models, they used GA or LBFGS to produce Sponge Examples. Sponge LBFGS samples perform better than GA on image classification models and can achieve around a maximum of 3% energy increase for processing compared to normal samples. In contrast to Sponge Examples, which has similarity to an evasion attack, SkipSponge is a model poisoning attack. Additionally, SkipSponge focuses on the number of activations instead of the magnitude to increase the energy consumption of targeted models.
Following this work, Cina et al. [10] introduced the Sponge Poisoning attack. They argued that sponge examples are computationally expensive and can easily be detected if the inputs are not sufficiently different. Moreover, they showed that using the number of non-zero activations ($\ell_{0}$ norm) in the objective increases the energy consumption more than using the magnitude of activations ($\ell_{2}$). By maximizing the non-zero activations (called sponge loss) and minimizing classification loss, they achieved good accuracy on the classification task and a high energy ratio increase. Sponge Poisoning was only performed on image classification models [10], but we also performed Sponge Poisoning on two GANs and two autoencoders, which required us to extend the ASIC simulator to support instance normalization layers and the Tanh activation function.
Sponge poisoning has also been applied to mobile phones. In particular, Paul et al. [39] found that Sponge Poisoning could increase the inference time on average by 13% and deplete the phone battery 15% faster on low-end devices. Wang et al. [50] showed that mobile devices that contain more advanced accelerators for deep learning computations may be more vulnerable to Sponge Poisoning.
Shapira et al.[42] were the first to consider sponging an object detection pipeline and focused on increasing the latency of various YOLO models. They achieved increased latency by creating a universal adversarial perturbation (UAP) on the input images with projected gradient descent with the L2 norm. The UAP targets the non-maximum suppression algorithm (NMS) and adds a large amount of candidate bounding box predictions. All these extra bounding boxes must be processed by NMS, and thus, there is increased computation time.
Finally, Hong et al. [16] introduced their weight poisoning attack and showed that an attacker could directly alter the values of weights in convolutional layers without significantly affecting the accuracy of a model. They used this attack to insert backdoors to deployed models. We build upon this idea of changing parameters directly to create SkipSponge. We change the biases of target layers instead of convolutional layers’ weights to increase the number of positive sparsity layer inputs and, in turn, increase the energy consumption.
@@
Shumailov et al. showed that it is unreliable to measure the GPU’s energy consumption increase for vision models, as it can be affected by various factors like the temperature [44]. Additionally, various ASIC accelerators with zero-skipping are discussed in previous sponge attacks [10, 44]. However, none of them [43, 14, 36, 1, 19] is implemented in silicon. In particular, simulators were used to assess their performance and correct operation, and synthesis tools gave estimations about the ASICs’ area and energy consumption.
VII Conclusions and Future Work
@@
This work proposes a novel sponge attack on DNNs. The SkipSponge attack changes the parameters of the pre-trained model. We show our attack is powerful (increasing energy consumption) and stealthy (making the detection more difficult). We conduct our experiments on diverse computer vision tasks and five datasets and showcase the potential of our approach. For future work, there are several interesting directions to follow. Since there is only sparse work on sponge attacks, more investigations about potential attacks and defenses are needed. Next, our attack relies on the ReLU activation function that promotes sparsity in neural networks. However, there are other (granted, much less used) activation functions that could potentially bring even more sparsity [25, 5]. Investigating the attack performance for those settings would be then interesting. Lastly, in this work, we used an ASIC accelerator simulator as provided by [44]. Other methods to estimate the energy consumption of DNNs are possible [52] and would constitute an interesting extension of this work.

### v5

@@ 5.8.2 Fine-pruning
CIFAR-10 95 $\rightarrow$ 94 9.6 $\rightarrow$ 6.3 88 $\rightarrow$ 89 7.1 $\rightarrow$ 6.4
@@
CIFAR-10 95 $\rightarrow$ 96 8.7 $\rightarrow$ 8.1 93 $\rightarrow$ 95 2.7 $\rightarrow$ 2.4
VGG-16 MNIST 94 $\rightarrow$ 98 11.8 $\rightarrow$ 11.9 97 $\rightarrow$ 98 8.9 $\rightarrow$ 7.8
CIFAR-10 86 $\rightarrow$ 90 4.0 $\rightarrow$ 1.9 89 $\rightarrow$ 75 32.6 $\rightarrow$ 31.6
@@
ResNet-18 MNIST 94 $\rightarrow$ 98 6.7 $\rightarrow$ 4.5 98 $\rightarrow$ 99 6.4 $\rightarrow$ 5.7
CIFAR-10 87 $\rightarrow$ 91 3.0 $\rightarrow$ 0.8 91 $\rightarrow$ 91 22.6 $\rightarrow$ 22.7
@@
5.8.3 Fine-tuning with Regularization
@@
Figure˜10 contains the results of the fine-tuning with regularization defense experiments. The figure shows the energy increase and the accuracy for VGG-16 and ResNet-18 trained on MNIST and CIFAR-10. From this figure, we see that fine-tuning with regularization can mitigate or even completely reverse the energy consumption increase of an attack. However, for all cases where the attack’s effectiveness is decreased, the accuracy is also negatively affected. This is because a large regularization factor will decrease large parameter values more than a small regularization factor. Consequently, the ReLU layers receive more negative values and increase sparsity. Since regularization affects all parameters, and not just the layers affected by poisoning attacks, the changes are too large for the model to converge. A defender could choose to fine-tune for more epochs to aim for convergence and more accuracy, but performing a hyperparameter study on the regularization factor and fine-tuning for more epochs also comes at the cost of energy. In Figure˜10, we see that for Sponge Poisoning and the CIFAR-10 dataset, the energy consumption is increased with a regularization factor of 1. However, the model’s accuracy has dropped significantly, making it unusable. The results of fine-tuning with regularization on other models and datasets show the same patterns as Figure˜10.
@@
Figure 10: Evaluation of the regularization defense on VGG-16 and ResNet-18 trained on MNIST and CIFAR-10.
5.9 Discussion
@@
We believe SkipSponge is a practical and important threat. It is effective against all tested models and is more effective than Sponge Poisoning [4] in increasing the GANs’ and the autoencoders’ energy consumption. Although SkipSponge results in a smaller energy increase than Sponge Poisoning for some image classification models, it cannot be easily spotted by a defender through an analysis of the activations of the model. Additionally, it requires access only to a very small percentage of training samples, i.e., one batch of samples may be enough, and the model’s weights. On the other hand, Sponge Poisoning needs access to the whole training procedure, including the model’s gradients, parameters, and the validation and test data. Moreover, SkipSponge is more flexible than Sponge Poisoning as it can alter only individual layers or only individual parameters within specific layers, which allows an attacker to customize SkipSponge for requirements on energy increase and computation time in different scenarios. Sponge Poisoning alters the entire model. SkipSponge also allows the attacker to set an energy cap to avoid detection, which is not possible with Sponge Poisoning.
6 Related Work
@@
The first sponge attack, called Sponge Examples, was introduced by Shumailov et al. [2]. Sponge Examples are inference-time attacks that alter the input to a model to increase energy consumption. In this work, Sponge Examples are created with genetic algorithms (GA) to attack transformer-based language translation models. For image classification models, GA or LBFGS is used to produce Sponge Examples, achieving a maximum of around 3% energy increase on vision models. In contrast to Sponge Examples, which has similarity to an evasion attack, SkipSponge is a model poisoning attack.
@@
Shumailov et al. [2] also showed that it is unreliable to directly measure a real GPU’s energy consumption increase for vision models, as it can be affected by various factors like temperature. To get around this issue, they proposed an ASIC simulator which we utilized in this work and discussed in Section˜2.3. Additionally, various ASIC accelerators with zero-skipping are discussed in previous sponge attacks [4, 2]. However, none of them [46, 22, 16, 21, 47] are implemented in silicon. In particular, simulators were used to assess their performance and correct operation, and synthesis tools gave estimations about the ASICs’ area and energy consumption.
Following Sponge Examples, Cina et al. [4] introduced the Sponge Poisoning attack. By altering the model’s training procedure to maximize the non-zero activations (called sponge loss) and minimize classification loss, they achieved good accuracy on the classification task and a high energy ratio increase. Sponge Poisoning was only performed on image classification models [4]. In our work, we performed Sponge Poisoning on two GANs and two autoencoders, which required us to extend the ASIC simulator to support instance normalization layers and the Tanh activation function.
Sponge Poisoning has also been applied to mobile phones. In particular, Paul et al. [5] found that Sponge Poisoning could increase the inference time on average by 13% and deplete the phone battery 15% faster on low-end devices.
Shapira et al.[3] were the first to consider sponge attacks against object detection models and focused on increasing the latency of the YOLO architecture. They increased the latency by creating a universal adversarial perturbation (UAP) on the input images with projected gradient descent with the L2 norm. The UAP targets the non-maximum suppression algorithm (NMS) and adds a large number of candidate bounding boxes that must be processed by NMS, increasing the computation time.
Finally, Hong et al. [27] showed that an attacker could directly alter the values of weights in convolutional layers without significantly affecting the accuracy of a model. They used this finding to insert backdoors into deployed models. We build upon this idea to create SkipSponge. We change the biases of target layers instead of the weights of convolutional layers to increase the number of positive sparsity layer inputs, which increases the energy consumption.
7 Conclusions and Future Work
@@
This work proposes a novel sponge attack on DNNs. The SkipSponge attack changes the parameters of the pretrained model. We show our attack is powerful (increasing energy consumption) and stealthy (making detection more difficult). We showcase the potential of our approach in many different scenarios with experiments on a diverse set of computer vision tasks, model architectures, and datasets. For future work, there are several interesting directions to follow. Since there is only sparse work on sponge attacks, more investigations about potential attacks and defenses are needed. Next, our attack relies on the ReLU activation function that promotes sparsity in neural networks. However, there are other (granted, much less used) activation functions that could potentially bring even more sparsity [48, 49]. Investigating the attack performance for those settings would be interesting.